OpenStudy (anonymous):

a rectanglr has a width of sqrt 63 and a length of sqrt 7+8. Find the area and the perimeter. express in simpliest form

OpenStudy (anonymous):

Perimeter \[2\sqrt{63} + 2\sqrt{7} + 16 = 6\sqrt{7} + 2\sqrt{7} + 16= 8\sqrt{7} + 16\] Area \[(\sqrt{63})(\sqrt{7} + 8) = (3\sqrt{7})(\sqrt{7} + 8)= 21 + 24\sqrt{7}\]